OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3)

OpenStudy (souvik):

place the terms with x in one side

OpenStudy (anonymous):

Ok Thanks!!

OpenStudy (anonymous):

make sure you don't multiply by any negative sign,or else the inequality gets changed.

OpenStudy (anonymous):

Oh ok! Thanks but can you help me step by step because I am kind of new to this.

OpenStudy (souvik):

\[-2(x-3)\ge 5-(x+3)\] or\[(-2)*x+(-2)*(-3)\ge 5-x-3\] or\[-2x+6\ge 2-x\] or\[6-2\ge -x+2x\] or\[4\ge x\] or \[x \le 4\]

OpenStudy (anonymous):

Thank You @souvik !!!!!!! It really helped me understand it more!!!

OpenStudy (souvik):

you are welcome....:)